The vapour pressure of aq.sol. of glucose is 750 mm of Hg at 373 K. Mole fraction is called as 
    (A)    0.087
    (B)    0.0087
    (C)    0.87
    (D)    0.013

Grade:upto college level

1 Answers

India VK
10 years ago

Lowering of VP = 760 - 750 = 10 mm Hg

 

So mole fraction of solute = 10/760 = 0.013